How to solve - 5/6b + 3/4 = 1/2b - 2/3 ( those are fractions
Answer by rfer(16322)

-5/6b-1/2b=-2/3-3/4
-5/6b-3/6b=-8/12-9/12
-8/6b=-17/12
b=(-17/12)/(-8/6)
b=-17/12*-6/8
b=102/96
b=1 1/16